OpenStudy (anonymous):

y=3-x find slope and yintercept?

jimthompson5910 (jim_thompson5910):

y=3-x y = -x + 3 y = -1x + 3

OpenStudy (anonymous):

i got -1 for slope and 3 for y intercept is this correct?

jimthompson5910 (jim_thompson5910):

compare this to y = mx+b form

jimthompson5910 (jim_thompson5910):

good, m = -1 is the slope b = 3 is the y-intercept, which is the point (0,3)

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np